# Dynamical systems defined by polynomials with algebraic properties

### v4

@@ Abstract
Let $St(\mathbb{T})$ be the set of streams, namely the two side sequences whose terms take value from the torus $\mathbb{T}$. We are mainly interested in the algebraic and dynamical properties of the space of stream zeros $\Omega_{P}$, which is the kernel via convolution product by a primitive polynomial $P(z)\in\mathbb{Z}[z]$ acting on $St(\mathbb{T})$. Together with the shift $\sigma$, it defines a topological dynamics $(\Omega_{P},\sigma)$ whose symbolic representation is given by the integer part of the action. It is also proved that two primitive polynomials $P(z),Q(z)$ are coprime if and only if $\dim(\Omega_{P}\cap\Omega_{Q})=0$. Furthermore, the group of (strong) automorphisms of $\Omega_{P}$ commuting with the shift $\sigma$ is also investigated. Specially in the case of $\deg P=2$, it is determined by the discriminant of equation $P(z)=0$. Finally, we study the streams over the Galois field $GF(q^{k})=GF(q)(\theta)$, where $\theta$ is a root of an irreducible polynomial $P(z)\in GF(q)[z]$ of degree $k$. The zeros $\Omega_{P}^{q}$ of $ST(GF(q^{k}))$ by the action of $P(z)$ in the above sense is discussed. In this case, the group of automorphisms of $\Omega_{P}^{q}$ commuting with the shift is proved to be a cyclic group of order $q^{k}-1$.
@@ 1 Introduction
Let $A=(a_{n};n\in{Z})$ be a two-side sequence of elements in an additive group $G$, which is called a two side stream over $G$. The set of two side streams over $G$ is denoted by $St(G)$. This $A$ is written as $(\cdots,a_{-2},a_{-1};a_{0},a_{1},\cdots)$ and is also viewed as a two-sided formal series on a variable,say $z$, namely,
$A=\cdots+a_{-2}z^{-2}+a_{-1}z^{-1}+a_{0}+a_{1}z+a_{2}z^{2}+\cdots.$
Note that $St(G)$ is an additive group by component-wise addition. Moreover, if $G$ is a topological group, then $St(G)$ is considered as a topological group with the product topology coming from $G$. Let 00 be the additive identity of $G$, then the additive identity of $St(G)$ is $(0^{\infty};0^{\infty})$, where $0^{\infty}$ before “;” implies $\cdots,0,0$ while that after “;” implies $0,0,\cdots$.
Let $G$ be a topological commutative ring having the multiplication identity 1. We define the convolution product of $(a_{n};n\in\mathbb{Z})$ and $(b_{n};n\in\mathbb{Z})\in St(G),$ if the following infinite series exists:
$(a_{n};n\in\mathbb{Z})\times(b_{n};n\in\mathbb{Z})=\left(\sum_{i=-\infty}^{\infty}a_{i}b_{n-i};~{}n\in\mathbb{Z}\right).$
The identity of $St(G)$ with respect to the convolution product is $I=(0^{\infty};1,0^{\infty})$. We call $(a_{n};n\in\mathbb{Z})\in St(G)$ absolutely summable if the series $\sum_{n\in\mathbb{Z}}a_{n}$ converges to the same value that is independent of the order of the sum.
@@
For $(a_{n};n\in\mathbb{Z})\in St(G)$, even if $(b_{n};~{}n\in\mathbb{Z})\in St(G)$ such that
$(a_{n};n\in\mathbb{Z})\times(b_{n};~{}n\in\mathbb{Z})=I$ (1)
@@
An element $(a_{n};n\in\mathbb{Z})$ of $St(\mathbb{C})$ such that $a_{n}=0$ except for finitely many $n$ can be identified with a polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ admitting negative powers, where $S\subset\mathbb{Z}$ is the finite set of $n$ such that $a_{n}\neq 0$. We always assume that $P(z)$ is nontrivial, that is, $S\neq\emptyset$. We call $P(z)$ hyperbolic if $P(z)=0$ does not hold for any $z\in\mathbb{C}$ with $|z|=1$.
We are especially interested in polynomials with integer coefficients whose greatest common divisor (GCD) is 1. We admit polynomials with negative power, but in most cases, it does not lose generality to restrict them only with nonnegative powers, so we will do sometimes.
Let $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ be the torus, $P(z)$ acts on $St(\mathbb{T})$ by the convolution product in the following way
$(P(z),(x_{n};n\in\mathbb{Z}))\mapsto(a_{n};n\in S)\times(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}).$
@@
$\Omega_{P}:=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):P(z)\times(x_{n};n\in\mathbb{Z})=(0^{\infty};0^{\infty})\right\}.$
We show that $\Omega_{P}$ has a lot of similarities with the algebraic roots of $P(z)=0$. This might provide us with a way of understanding the roots of algebraic equations.
Let $GF(q)$ be the Galois field of size $q$, where $q$ is a power of a prime. Let $P(z)$ be a polynomial in $GF(q)[z]$. In the last section, we discuss the zeros of $P(z)$-action to $St(GF(q))$, which is denoted by $\Omega_{P}^{q}$.
@@ 2 Convolution product and its inverse
(i) Among $A,B,C\in St(\mathbb{C})$, two of them are absolutely summable and the remaining one is bounded, then the convolution products below exist and it holds that
$(A\times B)\times C=A\times(B\times C).$
(ii) For any absolutely summable $(a_{n};n\in\mathbb{Z})\in St(\mathbb{C})$, an absolutely summable $(b_{n};n\in\mathbb{Z})$ satisfying $(1)$ is unique if it exists. We call $(b_{n};n\in\mathbb{Z})$ as this convolution inverse of $(a_{n};n\in\mathbb{Z})$ and is denoted by $(a_{n};n\in\mathbb{Z})^{-1}$.
(iii) If both of $(a_{n};n\in\mathbb{Z}),~{}(b_{n};n\in\mathbb{Z})\in St(\mathbb{C})$ are absolutely summable, then so is their convolution product. Furthermore, if both of $(a_{n};n\in\mathbb{Z})^{-1},(b_{n};n\in\mathbb{Z})^{-1}$ exist, then
$\left((a_{n};n\in\mathbb{Z})\times(b_{n};n\in\mathbb{Z})\right)^{-1}$
exists as well and
$\left((a_{n};n\in\mathbb{Z})\times(b_{n};n\in\mathbb{Z})\right)^{-1}=(a_{n};n\in\mathbb{Z})^{-1}\times(b_{n};n\in\mathbb{Z})^{-1}$
holds. Moreover, we have
$\left((a_{n};n\in\mathbb{Z})^{-1}\right)^{-1}=(a_{n};n\in\mathbb{Z}).$
(iv) Any hyperbolic polynomial $P=(a_{n};n\in\mathbb{Z})\not\equiv 0$ ($P=(a_{n};n\in\mathbb{Z})\neq(0^{\infty};0^{\infty})$ which is better? has an (comment: ”a unique” seems to be better)absolutely summable inverse.
@@
(i) Let
$A=(a_{n};n\in\mathbb{Z}),~{}B=(b_{n};n\in\mathbb{Z}),~{}C=(c_{n};n\in\mathbb{Z}).$
Then for any $n\in\mathbb{Z}$, we prove that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|a_{i}b_{j}c_{k}|<\infty.$ (2)
By symmetry, we may assume that $A,B$ are absolutely summable and $C$ is bounded. Then, there exists a constant $L>0$ such that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|a_{i}b_{j}c_{k}|<L\sum_{i,j\in\mathbb{Z}}|a_{i}b_{j}|=L\sum_{i\in\mathbb{Z}}|a_{i}|\sum_{j\in\mathbb{Z}}|b_{j}|<\infty.$
Thus, $\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}a_{i}b_{j}c_{k}$ converges absolutely. Hence, we have
$\sum_{k\in\mathbb{Z}}\left(\sum_{i,j\in\mathbb{Z}\atop{i+j=n-k}}a_{i}b_{j}\right)c_{k}=\sum_{i\in\mathbb{Z}}a_{i}\left(\sum_{j,k\in\mathbb{Z}\atop{j+k=n-i}}b_{j}c_{k}\right)=\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}a_{i}b_{j}c_{k}$
and $(A\times B)\times C=A\times(B\times C)$.
(ii) Suppose that any of $A,B,C\in St(\mathbb{C})$ be absolutely summable and both of $B,C$ are convolution inverses of $A$. Then by (i), we have $B=B\times(A\times C)=(B\times A)\times C=C$. Thus, the absolutely summable convolution inverse of $A$ is unique.
@@
(iv) Let $P(z)=\sum_{n\in S}a_{n}z^{n}$ with a finite nonempty set $S=\{n\in\mathbb{Z};a_{n}\neq 0\}$. Consider the factorization of $P(z)$ and applying (ii), it is sufficient to prove that $P(z)$ being of the form, i.e., constant $C\neq 0$, $z,z^{-1}$ and $z-\omega$ with $0<|\omega|\neq 1$, has the absolutely summable inverse.
@@
Case 3: $P(z)=z-w$ with $|\omega|>1.$ Since
@@
Case 4: $P(z)=z-w$ with $|\omega|<1$. Since
@@
Thus, $P(z)$ has the absolutely summable inverse which is the convolution product of the above terms. ∎
@@ 3 Stream zeros and dynamical systems
Let $P(z)$ be a primitive polynomial. Then clearly $\Omega_{P}$ is a compact $\mathbb{Z}$-subgroup of $St(\mathbb{T})$. Moreover, the pair $(\Omega_{P},\sigma)$, where $\sigma$ is the shift, becomes a compact dynamical system, which is called the dynamical system introduced by $P(z)$. On the other hand, let $\{\Omega_{P}\}=(\{x_{n}\};n\in\mathbb{Z})\in St([0,1))$, where $\{x_{n}\}$ is the representative of $x_{n}\in\mathbb{T}$ in $[0,1)$. Then, $P(z)\times\omega\in St(K_{P})$ holds for any $\omega\in\{\Omega_{P}\}$, where $K_{P}$ is a finite subset of integers with
$K_{P}=\{k_{*},k_{*}+1,\cdots,k^{*}\}$
here
$k_{*}:=(\sum_{n\in S,a_{n}<0}a_{n}+1)\wedge 0\mbox{ and }k^{*}:=(\sum_{n\in S,a_{n}>0}a_{n}-1)\vee 0.$
Let $\Xi_{P}$ be the closure of set $P(z)\times\{\Omega_{P}\}$. Then $\Xi_{P}$ is a compact, shift-invariant subset of $K_{P}^{\mathbb{Z}}$. Together with the shift $\sigma$, we call $(\Xi_{P},\sigma)$ the symbolic representation of $(\Omega_{P},\sigma)$
@@
Let a primitive polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ be hyperbolic. Then, the dynamical systems $(\Omega_{P},\sigma)$ and $(\Xi_{P},\sigma)$ are conjugate. Hence, we have the following commutative diagram. The conjugate map ×$\times P(z)$ is almost continuous, that is, continuous except for a countable set.
Moreover, a necessary and sufficient condition for $\delta=(\delta(n);~{}n\in\mathbb{Z})\in St(\mathbb{Z})$ to be $\delta\in\Xi_{P}$ is that $P(z)^{-1}\times\delta\in St([0,1))$, which automatically implies that $P(z)^{-1}\times\delta\in\{\Omega_{P}\}$.
$\{\Omega_{P}\}$$\{\Omega_{P}\}$$\Xi_{P}$$\Xi_{P}$$\sigma$$\sigma$×$\times P(z)^{-1}$×$\times P(z)^{-1}$×$\times P(z)$×$\times P(z)$
@@
The first half of the statement is clear. We prove the second half. If $\delta\in\Xi_{P}$, then there exists $\omega\in\Omega_{P}$ such that $\delta=P(z)\times(\{\omega_{n}\};n\in\mathbb{Z})$. Therefore, we have
$P(z)^{-1}\times\delta=P(z)^{-1}\times P(z)\times(\{\omega_{n}\};n\in\mathbb{Z})=(\{\omega_{n}\};n\in\mathbb{Z})\in[0,1)^{\mathbb{Z}}.$
Conversely, assume that $\tau:=P(z)^{-1}\times\delta\in[0,1)^{\mathbb{Z}}$. Since $P(z)\times\tau=\delta\in K_{P}^{\mathbb{Z}}$, we have $P(z)\times\pi(\tau)=(0^{\infty};0^{\infty})$, and hence, $\pi(\tau)\in\Omega_{P}$. Here, $\pi:\mathbb{R}^{\infty}\to\mathbb{T}^{\infty}$ is the projection. Since $\tau=\{\pi(\tau)\}$, we have $\delta=P(z)\times\tau\in\Xi_{P}$. ∎
@@
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):~{}x_{n+2}=3x_{n+1}-x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\}\}.$
@@
$\Omega_{P}=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):~{}\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&3\end{array}\right)\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\end{array}\right)\mbox{(mod 1) for any }n\in\mathbb{Z}\right\}.$
@@
Similarly, we can handle the case $|\omega_{-}|=|\frac{3-\sqrt{5}}{2}|<1,$
$(z-\omega_{-})^{-1}=z^{-1}(1-\omega_{-}z^{-1})^{-1}=z^{-1}+\omega_{-}z^{-2}+\omega_{-}^{2}z^{-3}+\cdots,$
together with

### v5

@@ Abstract
Let $St(G)$ be the set of streams over an additive group $G$, that is, the set of two side sequences from $\prod_{-\infty}^{\infty}G$. We are mainly interested in the case that $G$ is the torus $\mathbb{T}=\mathbb{R}/\mathbb{Z}$. Given a primitive polynomial $P(z)\in\mathbb{Z}[z]$ and identifying $(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$ as the formal power series $\sum_{n\in\mathbb{Z}}x_{n}z^{n}$, we shall study algebraic and dynamical properties of the space of stream zeros $\Omega_{P}\subset St(\mathbb{T})$, which is the kernel of the action $P(z)$ on $St(\mathbb{T})$, where the action is the multiplication as formal power series over $z$, considering $X:=(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$ as the formal power series $\sum_{n\in\mathbb{Z}}x_{n}z^{n}$. Together with the shift $\sigma$ defined by $(\sigma X)_{n}=x_{n+1}~(\forall n\in\mathbb{Z})$, $\Omega_{P}$ becomes a compact dynamical system $(\Omega_{P},\sigma)$ whose symbolic representation is given by the integer part of the image of the action. We proved that there is a one-to-one correspondence between the factorization of $P(z)$ and the semi-direct sum decomposition of the additive group $\Omega_{P}$. Furthermore, the group of strong automorphisms of $\Omega_{P}$ is investigated. Specially in the case of $\deg P=2$, it is determined by the discriminant of $P(z)$ and can be represented by the continued fraction expansion of its roots. Finally, we study the streams over the Galois field $GF(q)$. In this case, the group of strong automorphisms of $\Omega_{P}^{q}$ is a cyclic group of order $q^{k}-1$, where $k$ is the degree of $P(z)$.
@@ 1 Introduction
Let $X=(x_{n}\in G;n\in\mathbb{Z})$ be a two-side sequence of elements in an additive group $G$, which is called a two side stream (or simply, stream) over $G$. The set of two side streams over $G$ is denoted by $St(G)$. This $X$ is written as $(\cdots,x_{-2},x_{-1};x_{0},x_{1},\cdots)$, here the semicolon “;” separates the coordinates $x_{n}~(n<0)$ from $x_{n}~(n\geq 0)$. It is considered also as a two-side formal power series in a variable $z$, namely,
$X=\cdots+x_{-2}z^{-2}+x_{-1}z^{-1}+x_{0}+x_{1}z+x_{2}z^{2}+\cdots.$
Throughout this paper, we use the variable $z$ in this sense.
$St(G)$ is considered as an additive group by component-wise addition. Moreover, if $G$ is a topological group, then $St(G)$ is considered as a topological group with the product topology coming from $G$. Let 0 be the additive unit of $G$, then the additive unit of $St(G)$ is $(0^{\infty};0^{\infty})$, where $0^{\infty}$ before “;” implies $\cdots,0,0$ while that after “;” implies $0,0,\cdots$.
Let $G$ be a topological commutative ring having the multiplication unit 1. The product of $(x_{n};n\in\mathbb{Z})$ and $(y_{n};n\in\mathbb{Z})\in St(G)$ as the formal power series is defined if it exists, and called convolution product as well, thta is
$(x_{n};n\in\mathbb{Z})\times(y_{n};n\in\mathbb{Z})=\left(\sum_{i=-\infty}^{\infty}x_{i}y_{n-i};~n\in\mathbb{Z}\right).$ (1)
The unit of $St(G)$ with respect to the convolution product is $I=(0^{\infty};1,0^{\infty})$. We call $(x_{n};n\in\mathbb{Z})\in St(G)$ absolutely summable if the series $\sum_{n\in\mathbb{Z}}x_{n}$ converges to the same value independent of the order of the sum.
@@
For $(x_{n};n\in\mathbb{Z})\in St(G)$, even if $(y_{n};~n\in\mathbb{Z})\in St(G)$ such that
$(x_{n};n\in\mathbb{Z})\times(y_{n};~n\in\mathbb{Z})=I$
@@
An element $(a_{n};n\in\mathbb{Z})$ of $St(\mathbb{C})$ such that $a_{n}=0$ except for finitely many $n$ can be identified with a polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ admitting negative powers, where $S\subset\mathbb{Z}$ is the finite set of $n$ such that $a_{n}\neq 0$.
Definition 1.
We call a polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}\in\mathbb{Z}[z]$ primitive if $S\subset\{0,1,\cdots,k\}$ with $k\geq 1$, $a_{0}\neq 0,a_{k}\neq 0$ and the greatest common divisor (GCD) of the coefficients $a_{0},a_{1},\cdots,a_{k}$ is $\pm 1$. We call $P(z)$ hyperbolic if $P(z)=0$ does not hold for any $z\in\mathbb{C}$ with $|z|=1$.
Let $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ be the torus, which is a topological additive group with $\mathbb{Z}$-action. Hence, $St(\mathbb{T})$ is considered as a topological additive group with $\mathbb{Z}$-action, where the $\mathbb{Z}$-action is coodinatewise. Let $P(z)\in\mathbb{Z}[z]$ be a primitive polynomial as in Definition 1. Then, $P(z)$ acts on $St(\mathbb{T})$ as the multiplication of formal power series or the convolution product (1). That is,
$(P(z),(x_{n};n\in\mathbb{Z}))\mapsto(a_{n};n\in S)\times(x_{n};n\in\mathbb{Z})\\ =(a_{k}x_{n-k}+\cdots+a_{1}x_{n-1}+a_{0}x_{n};n\in\mathbb{Z})\in St(\mathbb{T}).$
@@
$\Omega_{P}:=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});P(z)\times(x_{n};n\in\mathbb{Z})=(0^{\infty};0^{\infty})\right\}.$
We show that $\Omega_{P}$ has a lot of similarities with the algebraic roots of $P(z)=0$.
The paper is organized as follows. In Section 2 we show some basic properties of the convolution product which is useful for further sections. In Section 3 we discuss $\Omega_{P}$ from dynamical system viewpoint. In Section 4 we introduce the notion of lift of a steam zeros on $\mathbb{T}$, that is, a stream zeros on $\mathbb{R}$ whose projection becomes it and prove Theorem 3 which gives a way to get a lift if Condition (#) is satisfied. Moreover, we prove that any primitive $P(z)$ satisfies Condition (#), which is of independent interest. When the lifting matrix of size $l=1$, this was answered by M. Newman [3] and X. Zhan [11]. In Section 5 we use the tools from resultant of two polynomials, generalized Vandermonde’s determinant, and Hadamard product to establish one-to-one correspondence between the factorization of $P(z)$ and the semi-direct sum decomposition of the additive group $\Omega_{P}$. In Section 6, we turn to investigate the algebraic structure of $\Omega_{P}$, we prove that the group of strong automorphisms of $\Omega_{P}$ commuting with the shift can be represented by the matrices from $Saut_{P}$ (Definition 4). In Section 7, we discuss the case $\deg P(z)=2$ and show that $Saut_{P}$ is determined by the discriminant of $P(z)$, while when the degree is larger than 2, we obtain only partial results, which also indicates that it has close relationship with the theory of Diophantine equations, known as Pell’s equation shown in the former case. We provide some examples to illustrate all theses theorem. In the last section, we focus on the zeros of $P(z)$ action on $St(GF(q))$, where $GF(q)$ is the Galois field of size $q$. We prove that if $P(z)$ is irreducible, then $Saut_{P}$ is a cyclic group of order $q^{k}-1$.
@@ 2 Convolution product and its inverse
Let $U=(u_{n};n\in\mathbb{Z}),~V=(v_{n};n\in\mathbb{Z}),~W=(w_{n};n\in\mathbb{Z})$ be arbitrary elements in $St(\mathbb{C})$.
(i) We have $U\times V=V\times U$ if one of $U\times V$ or $V\times U$ exists. If among $U,V,W$, two of them are absolutely summable and the remaining one is bounded, then the convolution products below exist and it holds that
$(U\times V)\times W=U\times(V\times W).$
(ii) If $U$ is absolutely summable, then an absolutely summable $V$ satisfying $U\times V=I$ or $V\times U=I$ is unique if exists. In this case, we have $U\times V=V\times U=I$ and We call $V$ as this convolution inverse of $U$ and is denoted by $U^{-1}$.
(iii) If both of $U,V\in St(\mathbb{C})$ are absolutely summable, then so is $U\times V$. Furthermore, if both of $U^{-1},V^{-1}$ exist, then $(U\times V)^{-1}$ exists and $(U\times V)^{-1}=V^{-1}\times U^{-1}$ holds. Moreover, $(U^{-1})^{-1}=U$ holds.
(iv) Any primitive polynomial $P(z)=(a_{n};n\in\mathbb{Z})$ has a unique absolutely summable inverse $P(z)^{-1}$ in $St(\mathbb{R})$ if it is hyperbolic.
@@
(i) The firsr statement is clear. For the second statement, let $U,V,W$ satisfy the required condition. Then for any $n\in\mathbb{Z}$, we prove that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|u_{i}v_{j}w_{k}|<\infty.$
By symmetry, we may assume that $U,V$ are absolutely summable and $W$ is bounded. Then, there exists a constant $L>0$ such that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|u_{i}v_{j}w_{k}|<L\sum_{i,j\in\mathbb{Z}}|u_{i}v_{j}|=L\sum_{i\in\mathbb{Z}}|u_{i}|\sum_{j\in\mathbb{Z}}|v_{j}|<\infty.$
Thus, $\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}u_{i}v_{j}w_{k}$ converges absolutely. Hence, we have
$\sum_{k\in\mathbb{Z}}\left(\sum_{i,j\in\mathbb{Z}\atop{i+j=n-k}}u_{i}v_{j}\right)w_{k}=\sum_{i\in\mathbb{Z}}u_{i}\left(\sum_{j,k\in\mathbb{Z}\atop{j+k=n-i}}v_{j}w_{k}\right)=\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}u_{i}v_{j}w_{k}$
and $(U\times V)\times W=U\times(V\times W)$.
(ii) We prove the uniqueness since the other statements follow from the first statement in (i). If $U\times V=U\times V^{\prime}=I$, then by (i), we have
$V^{\prime}=(UV)V^{\prime}=(VU)V^{\prime}=V(UV^{\prime})=V.$
@@
(iv) Let $P(z)=\sum_{n\in S}a_{n}z^{n}$ with a finite nonempty set $S=\{n\in\mathbb{Z};a_{n}\neq 0\}$. Consider the factorization of $P(z)$ and applying (ii), it is sufficient to prove that $P(z)$ being of the form, either constant $C\neq 0$, $z,z^{-1}$ or $z-\omega$ with $0<|\omega|\neq 1$, has the absolutely summable inverse.
@@
Case 3: P(z)=zwP(z)=z-w with $|\omega|>1.$ Since
@@
Case 4: P(z)=zwP(z)=z-w with $|\omega|<1$. Since
@@
If $\omega\notin\mathbb{R}$ in Case 2 (or Case 3), then the conjugate $\overline{\omega}$ is also in Case 2 (or Case 3, respectively) and $(z-\omega)^{-1}(z-\overline{\omega})^{-1}\in St(\mathbb{R})$. Thus, $P(z)$ has the absolutely summable inverse $P(z)^{-1}\in St(\mathbb{R})$ as the product of the above terms. ∎
@@ 3 Stream zeros and dynamical systems
Let $P(z)$ be a primitive polynomial over $\mathbb{Z}$. Then clearly $\Omega_{P}$ is a compact $\mathbb{Z}$-subgroup of $St(\mathbb{T})$. Moreover, the pair $(\Omega_{P},\sigma)$, where $\sigma$ is the shift, becomes a compact dynamical system, which is called the dynamical system introduced by $P(z)$. On the other hand, let
$\{\Omega_{P}\}=\{(\{x_{n}\};n\in\mathbb{Z})\in St([0,1));~(x_{n};n\in\mathbb{Z})\in\Omega_{P}\},$
where $\{x_{n}\}$ is the representative of $x_{n}\in\mathbb{T}$ in $[0,1)$. Then, $P(z)\times X\in St(K_{P})$ holds for any $X\in\{\Omega_{P}\}$, where $K_{P}$ is a finite subset of integers with
$K_{P}=\{k_{*}+1,k_{*}+2,\cdots,k^{*}-1\},~\mbox{where}$
[…]
Let
$\Xi_{P}=\{P(z)\times\{\tilde{x}_{n}\};~\{\tilde{x}_{n}\}\in\{\Omega_{P}\}\}.$ (3)
Then $\Xi_{P}$ is a compact, shift-invariant subset of $K_{P}^{\mathbb{Z}}$. Together with the shift $\sigma$, we call $(\Xi_{P},\sigma)$ the symbolic representation of $(\Omega_{P},\sigma)$
@@
Let a primitive polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ be hyperbolic. Then, the dynamical systems $(\Omega_{P},\sigma)$ and $(\Xi_{P},\sigma)$ are conjugate. Hence, we have the following commutative diagram. The conjugate map $\times P(z)$ is almost continuous, that is, continuous except for a countable set.
Moreover, a necessary and sufficient condition for $\xi=(\xi(n);~n\in\mathbb{Z})\in St(\mathbb{Z})$ to be $\xi\in\Xi_{P}$ is that $P(z)^{-1}\times\xi\in St([0,1))$, which automatically implies that $P(z)^{-1}\times\xi\in\Omega_{P}$.
$\{\Omega_{P}\}$$\{\Omega_{P}\}$$\Xi_{P}$$\Xi_{P}$$\sigma$$\sigma$$\times P(z)^{-1}$$\times P(z)^{-1}$$\times P(z)$$\times P(z)$
@@
The first half of the statement is clear. We prove the second half. If $\xi\in\Xi_{P}$, then there exists $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ such that $\xi=P(z)\times(\{x_{n}\};n\in\mathbb{Z})$. Therefore, we have
$P(z)^{-1}\times\xi=P(z)^{-1}\times P(z)\times(\{x_{n}\};n\in\mathbb{Z})=(\{x_{n}\};n\in\mathbb{Z})\in[0,1)^{\mathbb{Z}}.$
Conversely, assume that $y:=P(z)^{-1}\times\xi\in[0,1)^{\mathbb{Z}}$ and $x=\pi(y)\in St(\mathbb{T})$, $\pi:\mathbb{R}\to\mathbb{T}$ being the projection. Since $P(z)\times y=\xi\in K_{P}^{\mathbb{Z}}\subset\mathbb{Z}^{\mathbb{Z}}$, we have $P(z)\times x=(0^{\infty};0^{\infty})\in St(\mathbb{T})$, and hence, $x\in\Omega_{P}$. Hence, $\xi=P(z)\times y=P(z)\times\{x\}$ for some $x\in\Omega_{P}$. Thus, $\xi\in\Xi_{P}$. ∎
@@
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~x_{n+2}=3x_{n+1}-x_{n}~\mbox{(mod 1) for any}~n\in\mathbb{Z}\}.$
@@
$\Omega_{P}=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&3\end{array}\right)\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\end{array}\right)\mbox{(mod 1) for any }n\in\mathbb{Z}\right\}.$
@@
Similarly since $|\omega_{-}|=|\frac{3-\sqrt{5}}{2}|<1$,
$(z-\omega_{-})^{-1}=z^{-1}(1-\omega_{-}z^{-1})^{-1}=z^{-1}+\omega_{-}z^{-2}+\omega_{-}^{2}z^{-3}+\cdots.$
Since